OpenStudy (anonymous):

You have an open question: The area of a rectangle is 7 - 2/5 square feet. The length of this rectangle is 3 - 1/12 feet. what is the width?

OpenStudy (anonymous):

you know it

OpenStudy (anonymous):

u know this one?

OpenStudy (anonymous):

ill fan yall for right answer please

OpenStudy (anonymous):

hilbertboy96 u know it

OpenStudy (anonymous):

well this is simple to be honest. Draw a picture first.  The area of a rectangle is Area=length x width. So to obtain the width just do Area/length = width so \[\frac{ 7\frac{ 2 }{ 5 } }{ 3\frac{ 1 }{ 12 } }\]